OpenStudy (anonymous):

How would you compute the indefinite integral of 8cos(3x)sin(x) using complex exponentials?

OpenStudy (anonymous):

\[ \cos(3 x)= \frac {e^{3ix}+e^{-3ix}}{2}\\ \sin (x) = \frac {e^{ix}-e^{-ix}}{2i}\\ \]

OpenStudy (anonymous):

Now multiply out and integrate

OpenStudy (anonymous):

Oh okay! So you just have to substitute those in, got it thanks.

OpenStudy (anonymous):

Yes